OpenStudy (anonymous):

find which of the following expressions is a factor of y^7-128

OpenStudy (anonymous):

\[y-2\] is one factor

OpenStudy (anonymous):

but how do i know...can you show me some work?? thank you

OpenStudy (anonymous):

it helps to know that \[2^7=128\]

OpenStudy (anonymous):

then you have \[x^7-2^7\]

OpenStudy (anonymous):

Thank you Satellite73!

OpenStudy (anonymous):

and i guess you are supposed to know that \[a^7-b^7=(a-b)(a^6+a^5b+a^4b^2+a^3b^3+a^2b^4+ab^5+b^6)\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

THANKS A LOT!